find the range of each function for the domain {-1,0.5,3.7}
y=|x|-1

Answer by rapaljer(4551)

To find the range, substitute each value of x into the formula and determine the y values:

y=|x|-1
x= -1, y = |-1| -1 =1-1 =0
x=0.5, y = |.5|-1 = .5 -1 = -.5
x=3.7, y = |3.7| -1 = 3.7-1= 2.7

Therefore the Range is { 0,-0.5, 2.7}
Because the answer is a set, be sure to use set braces, and the order does not matter.

R^2 at SCC